### v1

# BehaviorGPT: Smart Agent Simulation for Autonomous Driving with Next-Patch Prediction

Abstract
Simulating realistic interactions among traffic agents is crucial for efficiently validating the safety of autonomous driving systems. Existing leading simulators primarily use an encoder-decoder structure to encode the historical trajectories for future simulation. However, such a paradigm complicates the model architecture, and the manual separation of history and future trajectories leads to low data utilization. To address these challenges, we propose Behavior Generative Pre-trained Transformers (BehaviorGPT), a decoder-only, autoregressive architecture designed to simulate the sequential motion of multiple agents. Crucially, our approach discards the traditional separation between "history" and "future," treating each time step as the "current" one, resulting in a simpler, more parameter- and data-efficient design that scales seamlessly with data and computation. Additionally, we introduce the Next-Patch Prediction Paradigm (NP3), which enables models to reason at the patch level of trajectories and capture long-range spatial-temporal interactions. BehaviorGPT ranks first across several metrics on the Waymo Sim Agents Benchmark, demonstrating its exceptional performance in multi-agent and agent-map interactions. We outperformed state-of-the-art models with a realism score of 0.741 and improved the minADE metric to 1.540, with an approximately 91.6% reduction in model parameters.
Keywords: Multi-Agent Systems, Transformers, Decoder-Only Architecture, Autonomous Driving
1 Introduction
Autonomous driving has emerged as an unstoppable trend, with its rapid development increasing the demand for faithful evaluation of autonomy systems’ reliability [30]. While on-road testing can measure driving performance by allowing autonomous vehicles (AVs) to interact with the physical world directly, the high cost per test and the scarcity of safety-critical scenarios in the real world have hindered large-scale and comprehensive evaluation. As an alternative, validating system safety via simulation has become increasingly attractive [15, 46, 50, 42] as it enables rapid testing in diverse driving scenarios simulated at a low cost. This work focuses on smart agent simulation, i.e., simulating the behavior of traffic participants such as vehicles, pedestrians, and cyclists in the digital world, which is critical for efficiently validating and iterating behavioral policies for AVs.
A good simulator should be realistic, matching the real-world distribution of multi-agent behaviors to support the assessment of AVs’ ability to coexist with humans safely. To this end, researchers started by designing naive simulators, which mainly replay the driving logs collected in the real world [27, 29]. When testing new driving policies that deviate from the ones during data collection, agents in such simulators often exhibit unrealistic interactions with AVs, owing to the lack of reactivity to AVs’ behavior changes. To simulate reactive agents, traditional approaches [15, 28] apply traffic rules to control agents heuristically [43, 26], which may struggle to capture real-world complexity. Recently, the availability of large-scale driving data [6, 16, 48], the emergence of powerful deep learning tools [19, 45, 20], and the prosperity of related fields such as motion forecasting [17, 44, 54, 40, 53], have spurred the development of learning-based agent simulation [42, 4, 24, 49, 52] towards more precise matching of behavioral distribution. With the establishment of standard benchmarks like the Waymo Open Sim Agents Challenge (WOSAC) [30], which systematically evaluates the realism of agent simulation in terms of kinematics, map compliance, and multi-agent interaction, the research on data-driven simulation approaches has been further advanced [47, 33].
Existing learning-based agent simulators [42, 4, 24, 49, 52, 47, 33] mainly mirror the techniques from motion forecasting [17, 44, 54, 40, 53, 39, 18] and opt for an encoder-decoder architecture, presumably due to the similarity between the two fields. Typically, these models use an encoder to extract historical information and a decoder to predict agents’ future states leveraging the encoded features. This paradigm requires manually splitting the multi-agent time series into a history and a future segment, with the two segments being processed by separate encoders and decoders with heterogeneous architecture. For example, MVTA [47] constructs training samples by randomly selecting a “current” timestamp to divide sequences into history and future components. Others [49, 33] use fixed-length agent trajectories as historical scene context, conditioned on which the multi-agent future is sampled from the decoder. Nonetheless, the benefit of spending a large part of model parameters and computing budget on historical scene encoding is unclear. Using heterogeneous encoding and decoding modules also complicates the architecture, which may harm the scalability of models. Moreover, the manual separation of history and future leads to low utilization of data and computation: as every point in the sequence can be used for the separation, we believe a sample-efficient framework should be able to learn from every possible history-future pair from the sequence in parallel, which cannot be easily achieved by encoder-decoder solutions.
Inspired by the success of Large Language Models (LLMs) [35, 36, 5], we introduce the first decoder-only autoregressive architecture, termed Behavior Generative Pre-trained Transformers (BehaviorGPT), into the field of smart agent simulation to overcome the limitation of previous works. By applying homogeneous Transformer blocks [45] to the complete trajectory snippets without differentiating history and future, we arrive at a simpler, more parameter-efficient, and more sample-efficient behavior simulator with greater potential to scale up. Utilizing relative spacetime representations [53], BehaviorGPT symmetrically models each agent state in the sequence as if it were the “current” one and tasks each state with modeling subsequent states’ distribution during training. As a result, our framework maximizes the utilization of agent data for autoregressive modeling, avoiding wasting any learning signals available in the time series.
Autoregressive modeling with imitation learning, however, suffers from compounding errors [37] and causal confusion [12]. For the behavior simulation task, we observed that blindly mimicking LLMs’ training paradigm of next-token prediction [33], regardless of the difference in tokens’ semantics across tasks, will make these issues more prominent. For a next-token prediction model where each token embeds an agent state in seconds, achieving a low training loss can be simply copying and pasting the current token as the next one without performing any long-range interaction reasoning in space or time. To mitigate this problem, we introduce a Next-Patch Prediction Paradigm (NP3) that enables models to reason at the patch level of trajectories, as illustrated in Fig. 1. By enforcing models to autoregressively generate the next trajectory patch containing multiple time steps, which requires understanding the high-level semantics of agent behavior and capturing long-range spatial-temporal interaction, we prevent models from leveraging simple shortcuts during training. We equip BehaviorGPT with NP3 and attain superior performance on WOSAC [30] with merely 3M model parameters, demonstrating the effectiveness of our modeling framework for smart agent simulation.
The main contributions of this paper are as follows:
Proposed the first decoder-only autoregressive architecture for smart agent simulation, which consists of homogeneous Transformer blocks that can process complete agent sequences with high parameter and sample efficiency;
Developed the Next-Patch Prediction scheme to enhance models’ capability of long-range interaction reasoning, leading to more realistic multi-agent simulation over a long horizon;
Achieved the top-ranking result in the Waymo Open Sim Agents Challenge.
2 Related Work
2.1 Multi-Agent Traffic Simulation
Multi-agent traffic simulation is crucial for developing and testing autonomous driving systems. From early systems like ALVINN [34] to contemporary simulators such as CARLA [15] and SUMO [28], these platforms use heuristic driving policies to simulate agent reactive behaviors [8, 7, 11]. Recent innovations include realistic simulators like Nocturne and the integration of learned sim agents that employ complex dynamics models and varied input modalities. However, these approaches struggle to capture real-world complexity as policies based on simple heuristics are not robust enough to handle all possible scenarios and need frequent updates when encountering unsolvable situations. With the availability of large-scale data and deep learning approaches, the use of generative models like VAEs [42], GANs [24], Diffusion [52], and Autoregressive models [47, 33] has also been observed in the modeling of multi-agent trajectories, which greatly enhances the realism and predictive capabilities of these simulations. Given the temporal dependency of agents’ trajectory, Autoregressive models are naturally more suitable for this task, while other generative models require extra design to encode such dependencies. Among the existing autoregressive models, two representatives are MVTA [47] and Trajeglish [33], but their complicated scene context encoders make them have poor scalability and require more data to obtain acceptable performance. Hence, there is an urgent need to develop parameter- and sample-efficient model for multi-agent traffic simulation.
2.2 Patch in Transformer
The application of patches in Transformer models has demonstrated significant potential across various data modalities. PatchTST [32] introduces a patching mechanism that leverages channel-independence to retain local semantic information more effectively while significantly reducing computational complexity. In natural language processing, BERT [13] employs subword tokenization [38], while in computer vision, the Vision Transformer segments images into 16×16 patches [14]. This design has found similar applications in time series forecasting, aiming to retain local semantic information and reduce computational complexity. By dividing time series data into subseries-level patches, these patches serve as input tokens for the Transformer model, enabling the model to handle longer historical data and enhance forecasting accuracy. Moreover, patching has proven effective in self-supervised learning, significantly improving representation learning performance and achieving excellent fine-tuning results on large datasets [2, 22, 3]. As trajectory prediction also deals with time series data, the patching mechanism is expected to help models better capture the complicated interaction relationships among road users and promote prediction performance.
3 Methodology
This section presents the proposed BehaviorGPT for multi-agent behavior simulation, with Fig. 2 illustrating the overall framework. To begin with, we provide the formulation of our map-conditioned, multi-agent autoregressive modeling. Then, we detail the architecture of BehaviorGPT, which adopts a Transformer decoder with a triple-attention mechanism to operate sequences at the patch level. Finally, we present the objective for model training.
3.1 Problem Formulation
In multi-agent behavior simulation, we aim to simulate agents’ future behavior in dynamic and complex environments. Specifically, we define a scenario as the composite of a vector map and the states of agents over time steps. At each time step, the state of the -th agent includes the agent’s position, velocity, yaw angle, and bounding box size. The semantic type of agents (e.g., vehicles, pedestrians, and cyclists) are also available. Given the sequential nature of agent trajectories, we formulate the problem as sequential predictions over trajectory patches, where the prediction of a previous patch will affect that of the next patch. We define an agent-level trajectory patch as follows:
where is the number of time steps covered by a patch, indicates the number of patches, and represents the -th trajectory patch of the -th agent, with denoting the states within the patch. On top of , we use to denote the -th scene-level patch:
where incorporates all agents’ states at the -th patch. Next, we factorize the multi-agent joint distribution over scene-level patches according to the chain rule, following which we further factorize the distribution of each scene-level patch over the agents based on the assumption that each agent executes the plan independently within the horizon of a patch:
Here, we denote by the joint distribution of all agents’ states over all time steps conditioned on the map . Considering the multimodality of agents’ behavior over the horizon of a patch, we assume to be a mixture model consisting of modes:
where is the probability of the -th mode. Given the sequential nature of the states within a patch, Eq. 4 also conducts factorization over the states per mode using the chain rule.
3.2 Relative Spacetime Representation
In our autoregressive formulation, we treat each trajectory patch as the “current” patch that is responsible for estimating the next-patch distribution during training, contrasting many existing approaches that designate one current time step per sequence [49, 47, 23]. As a result, it is inefficient to employ the well-established agent- or polyline-centric representation from the field of motion forecasting [44, 54, 31, 40, 25, 51, 41], given that these representations are computed under the reference frames determined by one current time step per sequence. For this reason, we adopt the relative spacetime representation introduced in QCNet [53] to model the patches symmetrically in time, thereby allowing parallel next-patch predictions. Under this representation, the features of each map element and agent state are derived from coordinate-independent attributes (e.g., the semantic category of a map element and the speed of an agent state). On top of this, we effectively maintain the spatial-temporal relationships between input elements via relative positional embeddings. Specifically, we use and to index two different input elements and compute the relative spatial-temporal embedding by:
where is the relational embedding from to , is the Euclidean distance between them, is the angle between (the orientation of ) and (the displacement vector from to ), is the relative yaw angle from to , and is the time difference between them.
3.3 Map Tokenization and Agent Patching
Before performing spatial-temporal relational reasoning among the input elements of a traffic scenario, we must convert the raw information into high-dimensional embeddings. We first embed map information by sampling map points per meters from the road graph and tokenizing them according to their semantic category (e.g., lane centerlines, road edges, and crosswalks). The -th map point’s embedding is denoted by . On the other hand, we process agent states using attention-based patching to obtain patch-level embeddings of trajectories. For the -th agent’s state at time step , we employ an MLP to transform the speed, the velocity vector’s yaw angle relative to the bounding box’s heading, the size of the bounding box, and the semantic type of the agent, into a feature vector . To further acquire patch embeddings, we collect the feature vectors of consecutive agent states and apply the attention mechanism with relative positional embeddings to them:
where is the patch embedding of the -th agent at the -th patch, denotes the multi-head self-attention [45], denotes concatenation, and indicates the positional embedding of relative to , which is computed according to Eq. 5. Such an operation can be viewed as aggregating the features of into that of and using the embeddings fused with high-level semantics as the agent tokens in the subsequent modules.
3.4 Triple-Attention Transformer Decoder
After obtaining the map tokens and the patch embeddings of agents, we employ a Transformer decoder [45] with the triple-attention mechanism to model the spatial-temporal interactions among the scene elements. The triple-attention mechanism considers three distinct sources of relations in the scene: the temporal dependencies over the trajectory patches per agent, the regulations of the map elements on the agents, and the social interactions among agents, as illustrated in Figure 3.
Temporal Self-Attention. This module captures the relationships among the trajectory patches of each individual agent. Similar to decoder-only LLMs [35, 36, 5], this module leverages the multi-head self-attention (MHSA) with a causal mask to enforce each trajectory patch to only attend to its preceding patches of the same agent, which is compatible with our autoregressive formulation. We equip the temporal MHSA with relative positional embeddings, which is as shown below:
where and are the temporal-aware feature vector and the patch embedding for the -th agent at the -th patch, respectively, and embeds the relative position from to , which represents the spatial-temporal relationship between the patches and .
Agent-Map Cross-Attention. Unlike natural language, which only has a sequential dimension (similar to the time dimension in our task), we must also conduct spatial reasoning to consider the environmental influence on agents’ behavior. To facilitate the modeling of agent-map interactions, we apply the multi-head cross-attention (MHCA) to each trajectory patch in the scenario. Considering that a scenario may comprise an explosive number of map points and that an agent would not be influenced by map elements far away, we filter the key/value map elements in MHCA using the k-nearest neighbors algorithm. The agent-map cross-attention is formulated as:
where is the map-aware feature vector for the -th agent at the -th patch, is the embedding of the -th map point, is the relative positional embedding between the agent state and the -th map point, and denotes the k-nearest map neighbors of .
Agent-Agent Self-Attention. We further capture the social interactions among agents by applying the MHSA to the space dimension of the trajectory patches. In this module, we also utilize the locality assumption induced by the k-nearest neighbor selection for better computational and memory efficiency. Specifically, the map-aware features of trajectory patches are refined by:
where is the feature vector enriched with spatial interaction information among agents for the -th agent at the -th patch, contains the relative information between the -th and the -th agent at the -th patch, and filters the k-nearest agent neighbors of .
Overall Decoder Architecture. Each of the triple-attention modules is enhanced by commonly used components in Transformers [45], including feed-forward networks, residual connections [19], and Layer Normalization [1] in a pre-norm fashion. To enable higher-order relational reasoning, we interleave the three Transformer layers, each of which is applied twice. We denote the ultimate feature of the -th agent at the -th patch as , which will serve as the input of the prediction head for next-patch prediction modeling.
3.5 Next-Patch Prediction Head
Given the semantically rich and interaction-aware patch features output by the Transformer decoder, we develop a next-patch prediction head to model the multimodal distribution of agent trajectories. We employ an autoregressive masking mechanism during inference to sample agents’ trajectories patch by patch. Specifically, based on the attention output of a patch, the prediction head will estimate the distribution of the next patch. After sampling the next trajectory patch from the distribution, we concatenate the predicted patch with the original sequence and continue to reason about the subsequent patches. An example with a patch size of is shown in Figure 4.
The following describes the process of next-patch prediction regarding the -th patch of the -th agent. Given the attention output , we intend to estimate the parameters of the next patch’s mixture model, which we pre-define with modes. To begin with, we use an MLP to transform into , i.e., the mixing coefficient of the modes. On the other hand, we employ a GRU-based autoregressive RNN [10] to unroll the states within the next patch step by step, which respects the sequential nature of behavior simulation. Specifically, the hidden state of the RNN is initialized with and is updated by the newly predicted states step by step during the rollout. At each step of the RNN, we use an MLP to estimate the location and scale parameters of the next agent state’s distribution based on the hidden state. The whole process is formulated as follows:
where and are locations and scales. We consider the distribution of the next agent state as a multivariate marginal distribution, parameterizing the position and velocity components as Laplace distributions and the yaw angle as a Von Mises distribution.
3.6 Training Objective
To train BehaviorGPT, we employ the negative log-likelihood loss on the factorized as formulated previously:
Note that each ground-truth trajectory patch is transformed into the viewpoint of its previous patch. During training, we utilize teacher forcing to parallelize the modeling of next-patch prediction and lower the learning difficulty, but we do not use the ground-truth agent states when updating the RNN’s hidden states, intending to train the model to recover from its mistakes made in next-state prediction.
4 Experiments
We use the Waymo Open Motion Dataset (WOMD) [16] to evaluate the effectiveness of our BehaviorGPT. We first introduce the dataset and the evaluation metrics, then present the rollout results obtained by BehaviorGPT on the Waymo Sim Agents Benchmark. Finally, we conduct ablation study to further compare and analyze the performance of BehaviorGPT under various settings.
4.1 Dataset and Metrics
Our experiments are conducted on the WOMD v1.2.1. The dataset comprises 486,995/44,097/44,920 training/validation/testing scenarios. Each scenario includes -step historical observations and -step future observations, totaling seconds. We simulate up to agents and generate simulation steps ( seconds) per agent at -second intervals in an autoregressive and reactive manner for evaluation. Each agent requires simulations, each including x/y/z centroid coordinates and a heading value. We use NVIDIA RTX 4090 GPUs to train models, where using training data takes hours and using training data takes days.
We use various metrics to evaluate our method. The minADE measures the minimum average displacement error over multiple simulated trajectories, assessing trajectory accuracy. REALISM is the meta metric that expects the simulations to match real-world observations. LINEAR ACCEL and LINEAR SPEED evaluate the realism regarding acceleration and speed. ANG ACCEL and ANG SPEED measure angular acceleration and speed used in maneuvering. DIST TO OBJ assesses the distances to objects, while COLLISION and TTC evaluate collision frequency and time to collision. Finally, DIST TO ROAD and OFFROAD measure map compliance.
4.2 Results on Waymo Sim Agents Benchmark
We evaluated the performance of our model using the Waymo Sim Agents Benchmark [30] and reported the results in Table 1. Although there were minor changes in the official calculation of metrics such as LINEAR ACCEL and COLLISION in 2024, the minADE metric remains stable and consistent. Thus, we consider minADE to be the primary metric for comparison. Our performance surpasses the state-of-the-art Trajeglish [33] across multiple metrics while also outperforming several well-established motion prediction models, including Wayformer [31], MultiPath++ [44], and MTR [40]. Notably, BehaviorGPT achieves the lowest minADE, indicating superior trajectory accuracy. Its high realism score underscores the model’s ability to match the real-world distribution while also maintaining competitive performance in speed, acceleration, and collision metrics. These strengths indicate that BehaviorGPT is a robust and reliable choice for traffic agent simulation, showcasing its advanced capabilities in simulating agents’ behavior in dynamic environments. As of the time we submitted the paper, the Waymo Open Sim Agents Challenge 2024 had not ended. We will update the complete results on the leaderboard in the revised version of our paper.
Model | minADE () | REALISM | LINEAR ACCEL | LINEAR SPEED | ANG ACCEL | ANG SPEED | DIST TO OBJ | COLLISION | TTC | DIST TO ROAD | OFFROAD |
Random Agent [30] | 50.706 | 0.155 | 0.044 | 0.002 | 0.012 | 0.074 | 0.000 | 0.000 | 0.734 | 0.178 | 0.287 |
Wayformer [31] | 2.498 | 0.575 | 0.098 | 0.331 | 0.406 | 0.413 | 0.297 | 0.870 | 0.782 | 0.592 | 0.866 |
CAD [9] | 2.308 | 0.531 | 0.253 | 0.349 | 0.310 | 0.432 | 0.332 | 0.568 | 0.789 | 0.637 | 0.834 |
MultiPath++ [44] | 2.049 | 0.533 | 0.230 | 0.434 | 0.452 | 0.515 | 0.345 | 0.567 | 0.812 | 0.639 | 0.682 |
MVTA [47] | 1.866 | 0.636 | 0.220 | 0.439 | 0.480 | 0.533 | 0.374 | 0.875 | 0.829 | 0.654 | 0.893 |
MTR+++ [40] | 1.679 | 0.608 | 0.107 | 0.414 | 0.436 | 0.484 | 0.347 | 0.861 | 0.797 | 0.654 | 0.895 |
MVTE [47] | 1.674 | 0.645 | 0.222 | 0.445 | 0.481 | 0.535 | 0.383 | 0.893 | 0.832 | 0.664 | 0.908 |
Trajeglish [33] | 1.571 | 0.645 | 0.192 | 0.453 | 0.538 | 0.485 | 0.386 | 0.922 | 0.836 | 0.659 | 0.886 |
Waymo Baseline (2024) [30] | 7.514 | 0.398 | 0.116 | 0.043 | 0.439 | 0.252 | 0.215 | 0.390 | 0.755 | 0.480 | 0.442 |
MVTE (2024) [47] | 1.676 | 0.730 | 0.353 | 0.359 | 0.599 | 0.497 | 0.374 | 0.904 | 0.830 | 0.665 | 0.907 |
BehaviorGPT (Ours) | 1.540 | 0.741 | 0.315 | 0.355 | 0.530 | 0.469 | 0.377 | 0.952 | 0.828 | 0.666 | 0.930 |
Besides the benchmarking results, we further compare the parameter count of BehaviorGPT against several baselines. Table 2 demonstrates that BehaviorGPT, with only 3 million parameters, achieves a notable reduction in parameter count compared with state-of-the-art models like MTR++ and MVTA. When compared against Trajeglish, BehaviorGPT achieves an approximately 91.6% reduction in parameter count. Despite its minimal parameter count, BehaviorGPT outperforms these more complex models, as shown in Table 1. These results demonstrate the parameter efficiency of decoder-only architecture and also indicate that our approach can utilize the training samples more efficiently.
4.3 Impact of Hyperparameters
In BehaviorGPT, four model hyperparameters are expected to impact model performance: the patch size, the number of spatial neighbors, the number of modes, and the probability threshold for trajectory mode sampling. We trained and evaluated BehaviorGPT using different hyperparameters to investigate their impact on model performance in the following.
Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
BehaviorGPT (patch size=1) | 2.3752 | 0.6783 | 0.2088 | 0.4021 | 0.3201 | 0.8432 |
BehaviorGPT (patch size=5) | 1.5598 | 0.7272 | 0.3217 | 0.4622 | 0.3768 | 0.9077 |
BehaviorGPT (patch size=10) | 1.5203 | 0.7334 | 0.3022 | 0.4733 | 0.3796 | 0.9131 |
Table 3 presents the results of BehaviorGPT with varying patch sizes. By comparing the most representative metrics, it is evident that using a patch size of , meaning training and predicting with steps ( seconds), significantly outperforms the model without patching. Additionally, increasing the patch size to further enhances the performance, though the improvement is less substantial. These results demonstrate the benefits of incorporating the NP3 into agent simulation.
The next hyperparameter to be investigated is the number of spatial neighbors considered in Transformers. Table 4 shows that increasing the neighbor number from to can improve the performance on all metrics, which shows that our model can easily scale up with more computation.
Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
BehaviorGPT (k=16) | 1.6463 | 0.7207 | 0.2659 | 0.4510 | 0.6540 | 0.8926 |
BehaviorGPT (k=32) | 1.5930 | 0.7283 | 0.2691 | 0.4556 | 0.6550 | 0.9137 |
Lastly, the number of modes refers to the number of predicted futures in multi-modal distributions, where a larger number of modes may enhance the trajectory diversity. We also adopt top- sampling [21] to select the smallest set of most probable tokens whose cumulative probability exceeds , where a larger encourages more diverse outputs. Table 5 reveals that for the same top- value, increasing the number of modes () generally improves the realism and offroad metrics but has varied effects on other metrics. For the same , increasing achieves better realism and linear acceleration but slightly decreases minADE and other metrics.
Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
BehaviorGPT (=8, p=0.95) | 1.6132 | 0.7290 | 0.2630 | 0.4564 | 0.6520 | 0.9174 |
BehaviorGPT (=16, p=0.8) | 1.6278 | 0.7274 | 0.2652 | 0.4557 | 0.6547 | 0.9148 |
BehaviorGPT (=16, p=0.9) | 1.5671 | 0.7298 | 0.2791 | 0.4631 | 0.6606 | 0.9147 |
BehaviorGPT (=16, p=0.95) | 1.5479 | 0.7310 | 0.2883 | 0.4686 | 0.6625 | 0.9131 |
BehaviorGPT (=16, p=1.0) | 1.5276 | 0.7335 | 0.2959 | 0.4744 | 0.6630 | 0.9142 |
5 Conclusion
In this work, we introduced BehaviorGPT, a decoder-only autoregressive architecture designed to enhance smart agent simulation for autonomous driving. By applying homogeneous Transformer blocks to entire trajectory snippets and utilizing relative spacetime representations, BehaviorGPT simplifies the modeling process and maximizes data utilization. To enable high-level understanding and long-range interaction reasoning in space and time, we developed the Next-Patch Prediction Paradigm, which requires the model to predict patches of trajectories instead of short segments. Our results on the Waymo Open Sim Agents Challenge demonstrate that BehaviorGPT, with just 3 million parameters, achieves superior performance, highlighting its potential to improve the realism and scalability of autonomous vehicle simulations with significantly fewer parameters.

### v2

@@
 Abstract
-Simulating realistic interactions among traffic agents is crucial for efficiently validating the safety of autonomous driving systems.
-Existing leading simulators primarily use an encoder-decoder structure to encode the historical trajectories for future simulation.
-However, such a paradigm complicates the model architecture, and the manual separation of history and future trajectories leads to low data utilization.
-To address these challenges, we propose Behavior Generative Pre-trained Transformers (BehaviorGPT), a decoder-only, autoregressive architecture designed to simulate the sequential motion of multiple agents.
-Crucially, our approach discards the traditional separation between "history" and "future," treating each time step as the "current" one, resulting in a simpler, more parameter- and data-efficient design that scales seamlessly with data and computation.
-Additionally, we introduce the Next-Patch Prediction Paradigm (NP3), which enables models to reason at the patch level of trajectories and capture long-range spatial-temporal interactions.
-BehaviorGPT ranks first across several metrics on the Waymo Sim Agents Benchmark, demonstrating its exceptional performance in multi-agent and agent-map interactions.
-We outperformed state-of-the-art models with a realism score of 0.741 and improved the minADE metric to 1.540, with an approximately 91.6% reduction in model parameters.
-Keywords: Multi-Agent Systems, Transformers, Decoder-Only Architecture, Autonomous Driving
+Simulating realistic behaviors of traffic agents is pivotal for efficiently validating the safety of autonomous driving systems.
+Existing data-driven simulators primarily use an encoder-decoder architecture to encode the historical trajectories before decoding the future.
+However, the heterogeneity between encoders and decoders complicates the models, and the manual separation of historical and future trajectories leads to low data utilization.
+Given these limitations, we propose BehaviorGPT, a homogeneous and fully autoregressive Transformer designed to simulate the sequential behavior of multiple agents.
+Crucially, our approach discards the traditional separation between "history" and "future" by modeling each time step as the "current" one for motion generation, leading to a simpler, more parameter- and data-efficient agent simulator.
+We further introduce the Next-Patch Prediction Paradigm (NP3) to mitigate the negative effects of autoregressive modeling, in which models are trained to reason at the patch level of trajectories and capture long-range spatial-temporal interactions.
+Despite having merely 3M model parameters, BehaviorGPT won first place in the 2024 Waymo Open Sim Agents Challenge with a realism score of 0.7473 and a minADE score of 1.4147, demonstrating its exceptional performance in traffic agent simulation.
+Keywords: Multi-Agent Systems, Transformers, Generative Models, Autonomous Driving 11footnotetext: *Equal contribution.
 1 Introduction
 Autonomous driving has emerged as an unstoppable trend, with its rapid development increasing the demand for faithful evaluation of autonomy systems’ reliability [32].
-While on-road testing can measure driving performance by allowing autonomous vehicles (AVs) to interact with the physical world directly, the high cost per test and the scarcity of safety-critical scenarios in the real world have hindered large-scale and comprehensive evaluation.
+While on-road testing can measure driving performance by allowing autonomous vehicles (AVs) to interact with the physical world directly, the high testing cost and the scarcity of safety-critical scenarios in the real world have hindered large-scale and comprehensive evaluation.
 As an alternative, validating system safety via simulation has become increasingly attractive [14, 48, 53, 44] as it enables rapid testing in diverse driving scenarios simulated at a low cost.
@@
 A good simulator should be realistic, matching the real-world distribution of multi-agent behaviors to support the assessment of AVs’ ability to coexist with humans safely.
-To this end, researchers started by designing naive simulators, which mainly replay the driving logs collected in the real world [27, 29].
+To this end, researchers started by designing naive simulators that mainly replay the driving logs collected in the real world [27, 29].
 When testing new driving policies that deviate from the ones during data collection, agents in such simulators often exhibit unrealistic interactions with AVs, owing to the lack of reactivity to AVs’ behavior changes.
 To simulate reactive agents, traditional approaches [14, 28] apply traffic rules to control agents heuristically [45, 26], which may struggle to capture real-world complexity.
-Recently, the availability of large-scale driving data [6, 16, 48], the emergence of powerful deep learning tools [19, 45, 20], and the prosperity of related fields such as motion forecasting [17, 44, 54, 40, 53], have spurred the development of learning-based agent simulation [42, 4, 24, 49, 52] towards more precise matching of behavioral distribution.
+Recently, the availability of large-scale driving data [6, 15, 50], the emergence of powerful deep learning tools [19, 47, 20], and the prosperity of related fields such as motion forecasting [16, 46, 59, 42, 58], have spurred the development of data-driven agent simulation [44, 4, 24, 52, 56] towards more precise matching of behavioral distribution.
 With the establishment of standard benchmarks like the Waymo Open Sim Agents Challenge (WOSAC) [32], which systematically evaluates the realism of agent simulation in terms of kinematics, map compliance, and multi-agent interaction, the research on data-driven simulation approaches has been further advanced [49, 35].
@@
 Typically, these models use an encoder to extract historical information and a decoder to predict agents’ future states leveraging the encoded features.
-This paradigm requires manually splitting the multi-agent time series into a history and a future segment, with the two segments being processed by separate encoders and decoders with heterogeneous architecture.
-For example, MVTA [47] constructs training samples by randomly selecting a “current” timestamp to divide sequences into history and future components.
+This paradigm requires manually splitting the multi-agent time series into a historical and a future segment, with the two segments being processed by separate encoders and decoders with heterogeneous architecture.
+For example, MVTA [49] constructs training samples by randomly selecting a “current” timestamp to divide sequences into historical and future components.
 Others [52, 35] use fixed-length agent trajectories as historical scene context, conditioned on which the multi-agent future is sampled from the decoder.
-Nonetheless, the benefit of spending a large part of model parameters and computing budget on historical scene encoding is unclear.
-Using heterogeneous encoding and decoding modules also complicates the architecture, which may harm the scalability of models.
-Moreover, the manual separation of history and future leads to low utilization of data and computation: as every point in the sequence can be used for the separation, we believe a sample-efficient framework should be able to learn from every possible history-future pair from the sequence in parallel, which cannot be easily achieved by encoder-decoder solutions.
-Inspired by the success of Large Language Models (LLMs) [35, 36, 5], we introduce the first decoder-only autoregressive architecture, termed Behavior Generative Pre-trained Transformers (BehaviorGPT), into the field of smart agent simulation to overcome the limitation of previous works.
-By applying homogeneous Transformer blocks [45] to the complete trajectory snippets without differentiating history and future, we arrive at a simpler, more parameter-efficient, and more sample-efficient behavior simulator with greater potential to scale up.
+Nonetheless, the benefit of employing heterogeneous modules to separately encode the history and decode the future, at the cost of significantly complicating the architecture, is unclear.
+Moreover, the manual separation of history and future leads to low utilization of data and computation: as every point in the sequence can be used for the separation, we believe a sample-efficient framework should be able to learn from every possible history-future pair from the sequence in parallel, which cannot be easily achieved by encoder-decoder solutions owing to their heterogeneous processing for the historical and the future time steps.
+Inspired by the success of decoder-only Large Language Models (LLMs) [37, 38, 5], we introduce a fully autoregressive Transformer architecture, dubbed BehaviorGPT, into the field of smart agent simulation to overcome the limitations of previous works.
+By applying homogeneous Transformer blocks [47] to the complete trajectory snippets without differentiating history and future, we arrive at a simpler, more parameter-efficient, and more sample-efficient solution for agent simulation.
 Utilizing relative spacetime representations [58], BehaviorGPT symmetrically models each agent state in the sequence as if it were the “current” one and tasks each state with modeling subsequent states’ distribution during training.
-As a result, our framework maximizes the utilization of agent data for autoregressive modeling, avoiding wasting any learning signals available in the time series.
+As a result, our framework maximizes the utilization of traffic data for autoregressive modeling, avoiding wasting any learning signals available in the time series.
 Autoregressive modeling with imitation learning, however, suffers from compounding errors [39] and causal confusion [11].
-For the behavior simulation task, we observed that blindly mimicking LLMs’ training paradigm of next-token prediction [33], regardless of the difference in tokens’ semantics across tasks, will make these issues more prominent.
-For a next-token prediction model where each token embeds an agent state in seconds, achieving a low training loss can be simply copying and pasting the current token as the next one without performing any long-range interaction reasoning in space or time.
-To mitigate this problem, we introduce a Next-Patch Prediction Paradigm (NP3) that enables models to reason at the patch level of trajectories, as illustrated in Fig.
-1.
-By enforcing models to autoregressively generate the next trajectory patch containing multiple time steps, which requires understanding the high-level semantics of agent behavior and capturing long-range spatial-temporal interaction, we prevent models from leveraging simple shortcuts during training.
+Concerning the behavior simulation task, we observed that blindly mimicking LLMs’ training paradigm of next-token prediction [35], regardless of the difference in tokens’ semantics across tasks, will make these issues more prominent.
+For a next-token prediction model embedding tokens at Hz, a low training loss can be achieved by simply copying and pasting the current token as the next one without performing any long-range interaction reasoning in space or time.
+To mitigate this issue, we introduce the Next-Patch Prediction Paradigm (NP3) that enables models to reason at the patch level of trajectories, as illustrated in Figure 1.
+By enforcing models to autoregressively generate the next trajectory patch containing multiple time steps, which requires understanding the high-level semantics of agent behaviors and capturing long-range spatial-temporal interactions, we prevent models from leveraging trivial shortcuts during training.
 We equip BehaviorGPT with NP3 and attain superior performance on WOSAC [32] with merely 3M model parameters, demonstrating the effectiveness of our modeling framework for smart agent simulation.
-The main contributions of this paper are as follows:
-Proposed the first decoder-only autoregressive architecture for smart agent simulation, which consists of homogeneous Transformer blocks that can process complete agent sequences with high parameter and sample efficiency;
-Developed the Next-Patch Prediction scheme to enhance models’ capability of long-range interaction reasoning, leading to more realistic multi-agent simulation over a long horizon;
-Achieved the top-ranking result in the Waymo Open Sim Agents Challenge.
+Our main contributions are three-fold.
+First, we propose a fully autoregressive architecture for smart agent simulation, which consists of homogeneous Transformer blocks that process multi-agent long sequences with high parameter and sample efficiency.
+Second, we develop the Next-Patch Prediction scheme to enhance long-range interaction reasoning, leading to more realistic multi-agent simulation over a long horizon.
+Third, we achieve remarkable performance on the Waymo Open Motion Dataset, winning first place in the 2024 Waymo Open Sim Agents Challenge.
+[…]
 2 Related Work
 2.1 Multi-Agent Traffic Simulation
-Multi-agent traffic simulation is crucial for developing and testing autonomous driving systems.
-From early systems like ALVINN [34] to contemporary simulators such as CARLA [15] and SUMO [28], these platforms use heuristic driving policies to simulate agent reactive behaviors [8, 7, 11].
-Recent innovations include realistic simulators like Nocturne and the integration of learned sim agents that employ complex dynamics models and varied input modalities.
-However, these approaches struggle to capture real-world complexity as policies based on simple heuristics are not robust enough to handle all possible scenarios and need frequent updates when encountering unsolvable situations.
-With the availability of large-scale data and deep learning approaches, the use of generative models like VAEs [42], GANs [24], Diffusion [52], and Autoregressive models [47, 33] has also been observed in the modeling of multi-agent trajectories, which greatly enhances the realism and predictive capabilities of these simulations.
-Given the temporal dependency of agents’ trajectory, Autoregressive models are naturally more suitable for this task, while other generative models require extra design to encode such dependencies.
-Among the existing autoregressive models, two representatives are MVTA [47] and Trajeglish [33], but their complicated scene context encoders make them have poor scalability and require more data to obtain acceptable performance.
-Hence, there is an urgent need to develop parameter- and sample-efficient model for multi-agent traffic simulation.
-2.2 Patch in Transformer
+Multi-agent traffic simulation is essential for developing and testing autonomous driving systems.
+From early systems like ALVINN [36] to contemporary simulators such as CARLA [14] and SUMO [28], these platforms have used heuristic driving policies to simulate agents’ reactive behaviors [8, 7, 10].
+However, they struggle to capture real-world complexity since policies based on simple heuristics are not robust enough to handle all sorts of scenarios.
+With the availability of large-scale data and deep learning approaches, generative models like VAEs [44], GANs [24], Diffusion [56], and autoregressive models [49, 41, 35] have gained success in generating multi-agent motions, which greatly enhance the realism of simulations.
+Given the temporal dependency of agent trajectories, autoregressive models naturally fit the simulation task, while others require extra designs to capture such dependencies.
+Among the existing autoregressive models, two representatives are MotionLM [41] and Trajeglish [35].
+Both of them adopt an encoder-decoder paradigm, designing complicated scene context encoders to extract historical information before autoregressive decoding.
+In contrast, our approach is fully autoregressive similar to decoder-only LLMs [37, 38, 5], which eliminates the need for using heterogeneous modules to process the historical and future time steps and achieves higher efficiency in terms of data and parameters via simpler architectural design.
+2.2 Patching Operations in Transformers
 The application of patches in Transformer models has demonstrated significant potential across various data modalities.
-PatchTST [32] introduces a patching mechanism that leverages channel-independence to retain local semantic information more effectively while significantly reducing computational complexity.
-In natural language processing, BERT [13] employs subword tokenization [38], while in computer vision, the Vision Transformer segments images into 16×16 patches [14].
-This design has found similar applications in time series forecasting, aiming to retain local semantic information and reduce computational complexity.
-By dividing time series data into subseries-level patches, these patches serve as input tokens for the Transformer model, enabling the model to handle longer historical data and enhance forecasting accuracy.
-Moreover, patching has proven effective in self-supervised learning, significantly improving representation learning performance and achieving excellent fine-tuning results on large datasets [2, 22, 3].
-As trajectory prediction also deals with time series data, the patching mechanism is expected to help models better capture the complicated interaction relationships among road users and promote prediction performance.
+For instance, BERT [12] employs subword tokenization [40] for natural language processing, while ViT [13] segments images into 2D patches for visual understanding.
+The patching design has also found applications in time-series forecasting [51, 57, 34], aiming at retaining local semantics and reducing computational complexity [34].
+Moreover, it has shown the effectiveness in self-supervised learning, which has significantly facilitated representation learning and contributed to excellent fine-tuning results on large datasets [2, 21, 3].
+Since the task of agent simulation also involves time-series data, we expect the patching mechanism to help models effectively capture the spatial-temporal interactions in driving scenarios and enhance the realism of the generated motion.
+Our proposed Next-Patch Prediction Paradigm (NP3) utilizes patch-level tokens in autoregressive modeling and trains each token to generate the next patch that comprises multi-step motions, which shares some similarities to multi-token prediction in LLMs [17].
 3 Methodology
-This section presents the proposed BehaviorGPT for multi-agent behavior simulation, with Fig.
-2 illustrating the overall framework.
+This section presents the proposed BehaviorGPT for multi-agent behavior simulation, with Figure 2 illustrating the overall framework.
 To begin with, we provide the formulation of our map-conditioned, multi-agent autoregressive modeling.
@@
 3.1 Problem Formulation
-In multi-agent behavior simulation, we aim to simulate agents’ future behavior in dynamic and complex environments.
+In multi-agent traffic simulation, we aim to simulate agents’ future behavior in dynamic and complex environments.
 Specifically, we define a scenario as the composite of a vector map and the states of agents over time steps.
@@
 The semantic type of agents (e.g., vehicles, pedestrians, and cyclists) are also available.
-Given the sequential nature of agent trajectories, we formulate the problem as sequential predictions over trajectory patches, where the prediction of a previous patch will affect that of the next patch.
-We define an agent-level trajectory patch as follows:
+Given the sequential nature of agent trajectories, we formulate the problem as sequential predictions over trajectory patches, where the prediction of each patch will affect the subsequent patches.
+We define an agent-level trajectory patch as
 where is the number of time steps covered by a patch, indicates the number of patches, and represents the -th trajectory patch of the -th agent, with denoting the states within the patch.
-On top of , we use to denote the -th scene-level patch:
-where incorporates all agents’ states at the -th patch.
-Next, we factorize the multi-agent joint distribution over scene-level patches according to the chain rule, following which we further factorize the distribution of each scene-level patch over the agents based on the assumption that each agent executes the plan independently within the horizon of a patch:
-Here, we denote by the joint distribution of all agents’ states over all time steps conditioned on the map .
-Considering the multimodality of agents’ behavior over the horizon of a patch, we assume to be a mixture model consisting of modes:
-where is the probability of the -th mode. Given the sequential nature of the states within a patch, Eq. 4 also conducts factorization over the states per mode using the chain rule.
+On top of , we use to denote the -th multi-agent patch, where incorporates all agents’ states at the -th patch.
+Next, we factorize the multi-agent joint distribution over patches along the time axis according to the chain rule:
+where is the joint distribution of all agents’ states over all time steps conditioned on the map .
+Further, we factorize over agents the conditional distribution of multi-agent patches based on the assumption that agents plan their motions independently within the horizon of a patch:
+Considering the multimodality of agents’ behavior within the horizon of a patch, we assume to be a mixture model consisting of modes:
+where is the probability of the -th mode. Given the sequential nature of the states within a patch, we further conduct factorization over the states per mode using the chain rule:
+Such an autoregressive formulation can be interpreted as planning the patch-level behavior of each agent independently (Eq.
+(3)), freezing agents’ behavior mode per time steps (Eq.
+(4)), and autoregressively unrolling the next state under a specific behavior mode (Eq.
+(5)).
+Under this formulation, we can flexibly adjust the replan frequency during inference to control the reactivity of agents.
+For example, we can let agents execute steps of the planned motions and choose a new behavior mode after steps to react to the change in environments.
 3.2 Relative Spacetime Representation
@@
 As a result, it is inefficient to employ the well-established agent- or polyline-centric representation from the field of motion forecasting [46, 59, 33, 42, 25, 54, 43], given that these representations are computed under the reference frames determined by one current time step per sequence.
-For this reason, we adopt the relative spacetime representation introduced in QCNet [53] to model the patches symmetrically in time, thereby allowing parallel next-patch predictions.
-Under this representation, the features of each map element and agent state are derived from coordinate-independent attributes (e.g., the semantic category of a map element and the speed of an agent state).
+For this reason, we adopt the relative spacetime representation introduced in QCNet [58] to model the patches symmetrically in space and time, achieving simultaneous multi-agent prediction when implementing Eq.
+(3) and allowing parallel next-patch prediction for the modeling of Eq.
+(2).
+Under this representation, the features of each map element and agent state are derived from coordinate-independent attributes, e.g., the semantic category of a map element and the speed of an agent state.
 On top of this, we effectively maintain the spatial-temporal relationships between input elements via relative positional embeddings.
-Specifically, we use and to index two different input elements and compute the relative spatial-temporal embedding by:
-where is the relational embedding from to , is the Euclidean distance between them, is the angle between (the orientation of ) and (the displacement vector from to ), is the relative yaw angle from to , and is the time difference between them.
+Specifically, we use and to index two different input elements and compute the relative spatial-temporal embedding by
+where is the relational embedding from to , is the Euclidean distance between them, is the angle between (i.e., the orientation of ) and (i.e., the displacement vector from to ), / is the relative yaw/height from to , and is the time difference.
 3.3 Map Tokenization and Agent Patching
 Before performing spatial-temporal relational reasoning among the input elements of a traffic scenario, we must convert the raw information into high-dimensional embeddings.
-We first embed map information by sampling map points per meters from the road graph and tokenizing them according to their semantic category (e.g., lane centerlines, road edges, and crosswalks).
-The -th map point’s embedding is denoted by .
+We first embed map information by sampling points along map polylines every meters and tokenizing the semantic category of each -meter segment (e.g., lane centerlines, road edges, and crosswalks) via learnable embeddings.
+The -th polyline segment’s embedding is denoted by , which does not include any information about coordinates.
 On the other hand, we process agent states using attention-based patching to obtain patch-level embeddings of trajectories.
-For the -th agent’s state at time step , we employ an MLP to transform the speed, the velocity vector’s yaw angle relative to the bounding box’s heading, the size of the bounding box, and the semantic type of the agent, into a feature vector .
+For the -th agent’s state at time step , we employ an MLP to transform the speed, the velocity vector’s angle relative to the bounding box’s heading, the size of the bounding box, and the semantic type of the agent, into a feature vector .
 To further acquire patch embeddings, we collect the feature vectors of consecutive agent states and apply the attention mechanism with relative positional embeddings to them:
-where is the patch embedding of the -th agent at the -th patch, denotes the multi-head self-attention [45], denotes concatenation, and indicates the positional embedding of relative to , which is computed according to Eq.
-5.
-Such an operation can be viewed as aggregating the features of into that of and using the embeddings fused with high-level semantics as the agent tokens in the subsequent modules.
+where is the patch embedding of the -th agent at the -th patch, denotes the multi-head self-attention [47], denotes concatenation, and indicates the positional embedding of relative to computed according to Eq.
+(6).
+Such an operation can be viewed as aggregating the features of into and using the embeddings fused with high-level semantics as the agent tokens in the subsequent modules.
 3.4 Triple-Attention Transformer Decoder
-After obtaining the map tokens and the patch embeddings of agents, we employ a Transformer decoder [45] with the triple-attention mechanism to model the spatial-temporal interactions among the scene elements.
-The triple-attention mechanism considers three distinct sources of relations in the scene: the temporal dependencies over the trajectory patches per agent, the regulations of the map elements on the agents, and the social interactions among agents, as illustrated in Figure 3.
+After obtaining map tokens and the patch embeddings of agents, we employ a Transformer decoder [47] with the triple-attention mechanism to model the spatial-temporal interactions among scene elements.
+As illustrated in Figure 3, the triple-attention mechanism considers three distinct sources of relations in the scene, including the temporal dependencies over the trajectory patches per agent, the regulations of the map elements on the agents, and the social interactions among agents.
 Temporal Self-Attention.
 This module captures the relationships among the trajectory patches of each individual agent.
-Similar to decoder-only LLMs [35, 36, 5], this module leverages the multi-head self-attention (MHSA) with a causal mask to enforce each trajectory patch to only attend to its preceding patches of the same agent, which is compatible with our autoregressive formulation.
-We equip the temporal MHSA with relative positional embeddings, which is as shown below:
-where and are the temporal-aware feature vector and the patch embedding for the -th agent at the -th patch, respectively, and embeds the relative position from to , which represents the spatial-temporal relationship between the patches and .
+Similar to decoder-only LLMs [37, 38, 5], it leverages the multi-head self-attention (MHSA) with a causal mask to enforce each trajectory patch to only attend to the preceding patches of the same agent, accommodating our autoregressive formulation.
+The temporal MHSA is equipped with relative positional embeddings:
+where and are the temporal-aware feature vector and the patch embedding of the -th agent at the -th patch, respectively, and embeds the relative position from to , which represents the spatial-temporal relationship between the patches and .
 Agent-Map Cross-Attention.
-Unlike natural language, which only has a sequential dimension (similar to the time dimension in our task), we must also conduct spatial reasoning to consider the environmental influence on agents’ behavior.
+Unlike natural language which only has a sequence dimension, we must also conduct spatial reasoning to consider the environmental influence on agents’ behavior.
 To facilitate the modeling of agent-map interactions, we apply the multi-head cross-attention (MHCA) to each trajectory patch in the scenario.
-Considering that a scenario may comprise an explosive number of map points and that an agent would not be influenced by map elements far away, we filter the key/value map elements in MHCA using the k-nearest neighbors algorithm.
-The agent-map cross-attention is formulated as:
-where is the map-aware feature vector for the -th agent at the -th patch, is the embedding of the -th map point, is the relative positional embedding between the agent state and the -th map point, and denotes the k-nearest map neighbors of .
+Considering that a scenario may comprise an explosive number of map polyline segments and that an agent would not be influenced by map elements far away, we filter the key/value map elements in MHCA using the k-nearest neighbors algorithm [42, 54].
+The agent-map cross-attention is formulated as
+where is the map-aware feature vector for the -th agent at the -th patch, is the embedding of the -th map polyline segment, is the relative positional embedding between the agent state and the -th map polyline segment, and denotes the k-nearest map neighbors of .
 Agent-Agent Self-Attention.
@@
 In this module, we also utilize the locality assumption induced by the k-nearest neighbor selection for better computational and memory efficiency.
-Specifically, the map-aware features of trajectory patches are refined by:
+Specifically, the map-aware features of trajectory patches are refined by
 where is the feature vector enriched with spatial interaction information among agents for the -th agent at the -th patch, contains the relative information between the -th and the -th agent at the -th patch, and filters the k-nearest agent neighbors of .
 Overall Decoder Architecture.
-Each of the triple-attention modules is enhanced by commonly used components in Transformers [45], including feed-forward networks, residual connections [19], and Layer Normalization [1] in a pre-norm fashion.
-To enable higher-order relational reasoning, we interleave the three Transformer layers, each of which is applied twice.
+Each of the attention layers above is enhanced by commonly used components in Transformers [47], including feed-forward networks, residual connections [19], and Layer Normalization [1] in a pre-norm fashion.
+To enable higher-order relational reasoning, we stack multiple triple-attention blocks by interleaving the three Transformer layers.
 We denote the ultimate feature of the -th agent at the -th patch as , which will serve as the input of the prediction head for next-patch prediction modeling.
 3.5 Next-Patch Prediction Head
-Given the semantically rich and interaction-aware patch features output by the Transformer decoder, we develop a next-patch prediction head to model the multimodal distribution of agent trajectories.
-We employ an autoregressive masking mechanism during inference to sample agents’ trajectories patch by patch.
-Specifically, based on the attention output of a patch, the prediction head will estimate the distribution of the next patch.
-After sampling the next trajectory patch from the distribution, we concatenate the predicted patch with the original sequence and continue to reason about the subsequent patches.
-An example with a patch size of is shown in Figure 4.
+Given the interaction-aware patch features output by the Transformer decoder, we develop a next-patch prediction head to model the marginal multimodal distribution of agent trajectories, which estimates the distributional parameters of each patch’s successor.
 The following describes the process of next-patch prediction regarding the -th patch of the -th agent.
-Given the attention output , we intend to estimate the parameters of the next patch’s mixture model, which we pre-define with modes.
-To begin with, we use an MLP to transform into , i.e., the mixing coefficient of the modes.
-On the other hand, we employ a GRU-based autoregressive RNN [10] to unroll the states within the next patch step by step, which respects the sequential nature of behavior simulation.
-Specifically, the hidden state of the RNN is initialized with and is updated by the newly predicted states step by step during the rollout.
-At each step of the RNN, we use an MLP to estimate the location and scale parameters of the next agent state’s distribution based on the hidden state.
-The whole process is formulated as follows:
-where and are locations and scales.
-We consider the distribution of the next agent state as a multivariate marginal distribution, parameterizing the position and velocity components as Laplace distributions and the yaw angle as a Von Mises distribution.
+Based on the attention output , we intend to estimate the parameters of the next patch’s mixture model pre-defined with modes.
+First, we use an MLP to transform into , the mixing coefficient of the modes.
+In each mode, the conditional distribution of the next agent state, as depicted in Eq.
+(5), is considered a multivariate marginal distribution that parameterizes the position and velocity components as Laplace distributions and the yaw angle as a von Mises distribution.
+Based on this formulation, we employ a GRU-based autoregressive RNN [9] to unroll the states within the next patch step by step, with each step being conditioned on the previously predicted states.
+Specifically, The hidden state of the RNN is initialized with at for .
+At each step of the rollout, we use an MLP to estimate the location and scale parameters of the next agent state’s position and velocity based on the hidden state.
+On the other hand, the MLP also estimates the location and concentration parameters of the next yaw angle.
+The location parameters of the newly predicted state, including the 3D positions, the 2D velocities, and the yaw angle, are used to update the RNN’s hidden state directly without relying on the predicted scale/concentration parameters for sampling.
+The whole process is summarized as follows:
+where , , and are the location, scale, and concentration parameters in the -th mode, and is the -th learnable mode embedding.
 3.6 Training Objective
-To train BehaviorGPT, we employ the negative log-likelihood loss on the factorized as formulated previously:
+To train BehaviorGPT, we apply the negative log-likelihood loss to the factorized distribution of as formulated previously:
 Note that each ground-truth trajectory patch is transformed into the viewpoint of its previous patch.
-During training, we utilize teacher forcing to parallelize the modeling of next-patch prediction and lower the learning difficulty, but we do not use the ground-truth agent states when updating the RNN’s hidden states, intending to train the model to recover from its mistakes made in next-state prediction.
+During training, we utilize teacher forcing to parallelize the modeling of next-patch prediction and ease the learning difficulty, but we do not use the ground-truth agent states when updating the RNN’s hidden states, intending to train the model to recover from its mistakes made in next-state prediction.
 4 Experiments
-We use the Waymo Open Motion Dataset (WOMD) [16] to evaluate the effectiveness of our BehaviorGPT.
-We first introduce the dataset and the evaluation metrics, then present the rollout results obtained by BehaviorGPT on the Waymo Sim Agents Benchmark.
-Finally, we conduct ablation study to further compare and analyze the performance of BehaviorGPT under various settings.
+This section first introduces the dataset and the evaluation metrics used in our experiments, followed by presenting the implementation details and the rollout results obtained by BehaviorGPT on the Waymo Open Sim Agents Benchmark [32].
+Finally, we conduct ablation studies to further compare and analyze the performance of BehaviorGPT under various settings.
 4.1 Dataset and Metrics
-Our experiments are conducted on the WOMD v1.2.1.
+Our experiments are conducted on the Waymo Open Motion Dataset (WOMD) [15].
 The dataset comprises 486,995/44,097/44,920 training/validation/testing scenarios.
-Each scenario includes -step historical observations and -step future observations, totaling seconds.
-We simulate up to agents and generate simulation steps ( seconds) per agent at -second intervals in an autoregressive and reactive manner for evaluation.
-Each agent requires simulations, each including x/y/z centroid coordinates and a heading value.
-We use NVIDIA RTX 4090 GPUs to train models, where using training data takes hours and using training data takes days.
-We use various metrics to evaluate our method.
+Each scenario includes -step observations sampled at Hz, totaling seconds.
+Given -step initial states of the scenarios, we simulate up to agents and generate simulation steps per agent at -second intervals in an autoregressive and reactive manner.
+Each agent requires simulations comprising x/y/z centroid coordinates and a heading value.
+The results on the test set are obtained by utilizing the full training set, while the performance on the validation set is based on of training data unless specified.
+We use various metrics for evaluation.
 The minADE measures the minimum average displacement error over multiple simulated trajectories, assessing trajectory accuracy.
-REALISM is the meta metric that expects the simulations to match real-world observations.
-LINEAR ACCEL and LINEAR SPEED evaluate the realism regarding acceleration and speed.
-ANG ACCEL and ANG SPEED measure angular acceleration and speed used in maneuvering.
-DIST TO OBJ assesses the distances to objects, while COLLISION and TTC evaluate collision frequency and time to collision.
-Finally, DIST TO ROAD and OFFROAD measure map compliance.
-4.2 Results on Waymo Sim Agents Benchmark
-We evaluated the performance of our model using the Waymo Sim Agents Benchmark [30] and reported the results in Table 1.
-Although there were minor changes in the official calculation of metrics such as LINEAR ACCEL and COLLISION in 2024, the minADE metric remains stable and consistent.
-Thus, we consider minADE to be the primary metric for comparison.
-Our performance surpasses the state-of-the-art Trajeglish [33] across multiple metrics while also outperforming several well-established motion prediction models, including Wayformer [31], MultiPath++ [44], and MTR [40].
-Notably, BehaviorGPT achieves the lowest minADE, indicating superior trajectory accuracy.
-Its high realism score underscores the model’s ability to match the real-world distribution while also maintaining competitive performance in speed, acceleration, and collision metrics.
-These strengths indicate that BehaviorGPT is a robust and reliable choice for traffic agent simulation, showcasing its advanced capabilities in simulating agents’ behavior in dynamic environments.
-As of the time we submitted the paper, the Waymo Open Sim Agents Challenge 2024 had not ended.
-We will update the complete results on the leaderboard in the revised version of our paper.
-Model | minADE () | REALISM | LINEAR ACCEL | LINEAR SPEED | ANG ACCEL | ANG SPEED | DIST TO OBJ | COLLISION | TTC | DIST TO ROAD | OFFROAD |
-Random Agent [30] | 50.706 | 0.155 | 0.044 | 0.002 | 0.012 | 0.074 | 0.000 | 0.000 | 0.734 | 0.178 | 0.287 |
-Wayformer [31] | 2.498 | 0.575 | 0.098 | 0.331 | 0.406 | 0.413 | 0.297 | 0.870 | 0.782 | 0.592 | 0.866 |
-CAD [9] | 2.308 | 0.531 | 0.253 | 0.349 | 0.310 | 0.432 | 0.332 | 0.568 | 0.789 | 0.637 | 0.834 |
-MultiPath++ [44] | 2.049 | 0.533 | 0.230 | 0.434 | 0.452 | 0.515 | 0.345 | 0.567 | 0.812 | 0.639 | 0.682 |
-MVTA [47] | 1.866 | 0.636 | 0.220 | 0.439 | 0.480 | 0.533 | 0.374 | 0.875 | 0.829 | 0.654 | 0.893 |
-MTR+++ [40] | 1.679 | 0.608 | 0.107 | 0.414 | 0.436 | 0.484 | 0.347 | 0.861 | 0.797 | 0.654 | 0.895 |
-MVTE [47] | 1.674 | 0.645 | 0.222 | 0.445 | 0.481 | 0.535 | 0.383 | 0.893 | 0.832 | 0.664 | 0.908 |
-Trajeglish [33] | 1.571 | 0.645 | 0.192 | 0.453 | 0.538 | 0.485 | 0.386 | 0.922 | 0.836 | 0.659 | 0.886 |
-Waymo Baseline (2024) [30] | 7.514 | 0.398 | 0.116 | 0.043 | 0.439 | 0.252 | 0.215 | 0.390 | 0.755 | 0.480 | 0.442 |
-MVTE (2024) [47] | 1.676 | 0.730 | 0.353 | 0.359 | 0.599 | 0.497 | 0.374 | 0.904 | 0.830 | 0.665 | 0.907 |
-BehaviorGPT (Ours) | 1.540 | 0.741 | 0.315 | 0.355 | 0.530 | 0.469 | 0.377 | 0.952 | 0.828 | 0.666 | 0.930 |
-Besides the benchmarking results, we further compare the parameter count of BehaviorGPT against several baselines.
-Table 2 demonstrates that BehaviorGPT, with only 3 million parameters, achieves a notable reduction in parameter count compared with state-of-the-art models like MTR++ and MVTA.
-When compared against Trajeglish, BehaviorGPT achieves an approximately 91.6% reduction in parameter count.
-Despite its minimal parameter count, BehaviorGPT outperforms these more complex models, as shown in Table 1.
-These results demonstrate the parameter efficiency of decoder-only architecture and also indicate that our approach can utilize the training samples more efficiently.
-4.3 Impact of Hyperparameters
-In BehaviorGPT, four model hyperparameters are expected to impact model performance: the patch size, the number of spatial neighbors, the number of modes, and the probability threshold for trajectory mode sampling.
-We trained and evaluated BehaviorGPT using different hyperparameters to investigate their impact on model performance in the following.
-Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
-BehaviorGPT (patch size=1) | 2.3752 | 0.6783 | 0.2088 | 0.4021 | 0.3201 | 0.8432 |
-BehaviorGPT (patch size=5) | 1.5598 | 0.7272 | 0.3217 | 0.4622 | 0.3768 | 0.9077 |
-BehaviorGPT (patch size=10) | 1.5203 | 0.7334 | 0.3022 | 0.4733 | 0.3796 | 0.9131 |
+REALISM is the meta-metric that expects the simulations to match the real-world distribution.
+LINEAR SPEED and LINEAR ACCEL evaluate the realism regarding speed and acceleration.
+Similarly, ANG SPEED and ANG ACCEL measure the realism of angular speed and acceleration.
+DIST TO OBJ considers the distances to objects, while COLLISION and TTC assess the simulation performance in terms of collision and time to collision.
+Finally, DIST TO ROAD EDGE and OFFROAD focus on map compliance.
+4.2 Implementation Details
+The optimal patch size we experimented with is , corresponding to second.
+All hidden sizes are set to .
+Each attention layer has attention heads with dimensions per head.
+To save training resources, we limit the maximum number of agents per scenario to and restrict the maximum number of neighbors in kNN attention layers to .
+The prediction head produces modes per agent and time step.
+We train the models for epochs on NVIDIA RTX 4090 GPUs with a batch size of , utilizing the AdamW optimizer [31].
+The weight decay rate and dropout rate are both set to .
+The learning rate is initially set to and decayed to following a cosine annealing schedule [30].
+Our results in the 2024 WOSAC are obtained using a single model with decoding blocks and a total of 3M parameters.
+To produce replicas of rollouts, we randomly sample behavior modes from agents’ next-patch distributions until completing the -second multi-agent trajectories, and we repeat this process with different random seeds.
+The final results on the leaderboard are based on a replan rate of Hz, while the ablation studies are based on a -Hz replan rate unless specified.
+4.3 Quantitative Results
+| Model | #Param | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
+| Linear Extrapolation [32] | - | 7.5148 | 0.3985 | 0.0434 | 0.1661 | 0.2522 | 0.4393 | 0.2154 | 0.3905 | 0.7555 | 0.4801 | 0.4426 |
+| TrafficBotsV1.5 [55] | 10M | 1.8825 | 0.6988 | 0.3361 | 0.3497 | 0.4512 | 0.5844 | 0.3596 | 0.8083 | 0.8209 | 0.6423 | 0.9134 |
+| VBD [23] | 12M | 1.4743 | 0.7200 | 0.3591 | 0.3664 | 0.4197 | 0.5222 | 0.3683 | 0.9341 | 0.8153 | 0.6508 | 0.8788 |
+| MVTE [49] | >65M | 1.6770 | 0.7302 | 0.3506 | 0.3531 | 0.4974 | 0.6000 | 0.3743 | 0.9049 | 0.8310 | 0.6655 | 0.9071 |
+| GUMP [22] | 523M | 1.6041 | 0.7431 | 0.3567 | 0.4111 | 0.5089 | 0.6353 | 0.3707 | 0.9403 | 0.8276 | 0.6686 | 0.9028 |
+| BehaviorGPT (Ours) | 3M | 1.4147 | 0.7473 | 0.3615 | 0.3365 | 0.4806 | 0.5544 | 0.3834 | 0.9537 | 0.8308 | 0.6702 | 0.9349 |
+We report the test set results in Table 1.
+Notably, BehaviorGPT achieves the lowest minADE and the best REALISM, underscoring the model’s ability to match the real-world distribution.
+Its excellent performance on COLLISION and OFFROAD also indicates that the model has successfully captured the agent-agent and agent-map interactions in driving scenarios.
+Besides the benchmarking results, we also compare the number of model parameters in BehaviorGPT and other baselines.
+Table 1 demonstrates that BehaviorGPT, with only 3M parameters, achieves more realistic simulation than significantly larger models like MVTE [49] and GUMP [22], which demonstrates the parameter efficiency of our approach.
+Without employing tricks like data augmentation, model ensemble, or post-processing steps, BehaviorGPT won first place in the 2024 Waymo Open Sim Agents Challenge.
+4.4 Qualitative Results
+Figure 4 visualizes some qualitative results of the rollouts produced by our model.
+In this scenario, BehaviorGPT can generate multiple plausible futures given the same initial states of agents, which demonstrates its capability of simulating diverse yet realistic agent behavior.
+However, we also note that autoregressive models still suffer from accumulated errors in some cases.
+As shown in Figure 5, the vehicle in orange gradually goes out of the road as time goes by, which indicates the inherent limitations of autoregressive generation.
+4.5 Ablation Studies
+We conduct some ablation studies to gain a more in-depth understanding of our approach.
+Impact of patch size.
 Table 2 presents the results of BehaviorGPT with varying patch sizes.
-By comparing the most representative metrics, it is evident that using a patch size of , meaning training and predicting with steps ( seconds), significantly outperforms the model without patching.
-Additionally, increasing the patch size to further enhances the performance, though the improvement is less substantial.
+According to the results, it is evident that using a patch size of , i.e., training and predicting with -Hz tokens, significantly outperforms the baseline without patching.
+Moreover, increasing the patch size to further enhances the overall performance.
 These results demonstrate the benefits of incorporating the NP3 into agent simulation.
-The next hyperparameter to be investigated is the number of spatial neighbors considered in Transformers.
-Table 4 shows that increasing the neighbor number from to can improve the performance on all metrics, which shows that our model can easily scale up with more computation.
-Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
-BehaviorGPT (k=16) | 1.6463 | 0.7207 | 0.2659 | 0.4510 | 0.6540 | 0.8926 |
-BehaviorGPT (k=32) | 1.5930 | 0.7283 | 0.2691 | 0.4556 | 0.6550 | 0.9137 |
-Lastly, the number of modes refers to the number of predicted futures in multi-modal distributions, where a larger number of modes may enhance the trajectory diversity.
-We also adopt top- sampling [21] to select the smallest set of most probable tokens whose cumulative probability exceeds , where a larger encourages more diverse outputs.
-Table 5 reveals that for the same top- value, increasing the number of modes () generally improves the realism and offroad metrics but has varied effects on other metrics.
-For the same , increasing achieves better realism and linear acceleration but slightly decreases minADE and other metrics.
-Model Variants | minADE () | REALISM | LINEAR ACCEL | ANG SPEED | DIST TO OBJ | OFFROAD |
-BehaviorGPT (=8, p=0.95) | 1.6132 | 0.7290 | 0.2630 | 0.4564 | 0.6520 | 0.9174 |
-BehaviorGPT (=16, p=0.8) | 1.6278 | 0.7274 | 0.2652 | 0.4557 | 0.6547 | 0.9148 |
-BehaviorGPT (=16, p=0.9) | 1.5671 | 0.7298 | 0.2791 | 0.4631 | 0.6606 | 0.9147 |
-BehaviorGPT (=16, p=0.95) | 1.5479 | 0.7310 | 0.2883 | 0.4686 | 0.6625 | 0.9131 |
-BehaviorGPT (=16, p=1.0) | 1.5276 | 0.7335 | 0.2959 | 0.4744 | 0.6630 | 0.9142 |
+However, changing the patch size also leads to a variation in replan frequency, which also has an influence on simulation.
+Next, we investigate the impact of replan frequency on the test set using the model submitted to the 2024 WOSAC.
+| Patch Size | Replan Frequency | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
+| 1 | 10 Hz | 2.3752 | 0.6783 | 0.2559 | 0.2088 | 0.4022 | 0.5094 | 0.3201 | 0.9002 | 0.8015 | 0.6149 | 0.8432 |
+| 5 | 2 Hz | 1.5599 | 0.7273 | 0.3543 | 0.3218 | 0.4623 | 0.5435 | 0.3768 | 0.9181 | 0.8339 | 0.6564 | 0.9077 |
+| 10 | 1 Hz | 1.5203 | 0.7335 | 0.3517 | 0.3023 | 0.4734 | 0.5432 | 0.3797 | 0.9358 | 0.8329 | 0.6645 | 0.9132 |
+Impact of replan frequency.
+During inference, we vary the replan frequency of the model with a patch size of by discarding a portion of the predicted states at each simulation step.
+As shown in Table 3, increasing the replan frequency from Hz to Hz can even improve the overall performance, which may benefit from the enhanced reactivity.
+This phenomenon demonstrates that the performance gain is not merely due to the lower replan frequency, as the model with a patch size of beats that with a patch size of even harder if using the same replan frequency of Hz.
+However, using an overly high replan frequency harms the performance, as indicated by the third row of Table 3.
+Overall, we conclude that using a larger patch indeed helps long-term reasoning, but a moderate replan frequency is important for temporal stability, which may be neglected by prior works.
+| Patch Size | Replan Frequency | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
+| 10 | 1 Hz | 1.5405 | 0.7414 | 0.3553 | 0.3153 | 0.4695 | 0.5303 | 0.3772 | 0.9520 | 0.8285 | 0.6664 | 0.9308 |
+| 10 | 2 Hz | 1.4147 | 0.7473 | 0.3615 | 0.3365 | 0.4806 | 0.5544 | 0.3834 | 0.9537 | 0.8308 | 0.6702 | 0.9349 |
+| 10 | 5 Hz | 1.5693 | 0.7342 | 0.3430 | 0.3472 | 0.4663 | 0.5673 | 0.3722 | 0.9429 | 0.8253 | 0.6534 | 0.9089 |
+| Agent-Agent Self-Attention | minADE () | REALISM () | DIST TO OBJ () | COLLISION () | TTC () |
+| 2.1489 | 0.6659 | 0.3539 | 0.6987 | 0.8070 | |
+| ✓ | 1.6247 | 0.7349 | 0.3783 | 0.9409 | 0.8320 |
+Impact of multi-agent interaction modeling.
+We remove all agent-agent self-attention layers in the first row of Table 4 to show that modeling the interactions among agents can boost minADE and REALISM.
+In particular, the realism in terms of collision is improved by when employing agent-agent self-attention.
+| Train Data | #Param | minADE () | REALISM () |
+| 20% | 5M | 1.4881 | 0.7396 |
+| 50% | 5M | 1.4060 | 0.7427 |
+| 100% | 5M | 1.3804 | 0.7438 |
+| Model Depth | #Param | minADE () | REALISM () |
+| 2 | 3M | 1.6247 | 0.7349 |
+| 3 | 4M | 1.5381 | 0.7387 |
+| 4 | 5M | 1.4881 | 0.7396 |
+| Model Width | #Param | minADE () | REALISM () |
+| 64 | 800K | 1.9637 | 0.7251 |
+| 128 | 3M | 1.6247 | 0.7349 |
+| 192 | 7M | 1.4993 | 0.7382 |
+| Training | Inference | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
+| 9.1 sec | 9.1 sec | 1.6247 | 0.7349 | 0.3546 | 0.3105 | 0.4689 | 0.5363 | 0.3783 | 0.9409 | 0.8320 | 0.6605 | 0.9163 |
+| 5.0 sec | 9.1 sec | 1.6294 | 0.7333 | 0.3565 | 0.3471 | 0.4613 | 0.5293 | 0.3813 | 0.9375 | 0.8273 | 0.6585 | 0.9100 |
+Scaling with data.
+We train our models with different proportions of training data.
+All the models have decoding blocks and a hidden size of , totaling 5M parameters.
+As shown in Table 4.5, BehaviorGPT is able to achieve remarkable performance with merely of training data, which is attributed to the high data efficiency of our approach.
+Increasing the proportion of training data from to further improves the performance on minADE and REALISM, and training on of the data continues to gain enhancement.
+Judging from the trend in Table 4.5, we believe that feeding more data for model training will continuously achieve better simulation performance.
+Scaling with model size.
+We investigate the effects of scaling up the model size based on some preliminary experiments with of training data.
+In Table 4.5, we vary the number of decoding blocks while fixing the hidden size as .
+On the other hand, we fix the number of decoding blocks as and vary the hidden size, as depicted in Table 4.5.
+Based on the experimental results, we can summarize that enlarging the model consistently leads to more realistic simulation, which showcases the potential of BehaviorGPT for scaling up.
+Extrapolation ability.
+We tried training a model on -second sequences and generating -second sequences during inference.
+The results in Table 8 show that this model achieves similar performance compared with the baseline trained with -second sequences, demonstrating our approach’s extrapolation ability to generate longer sequences.
 5 Conclusion
-In this work, we introduced BehaviorGPT, a decoder-only autoregressive architecture designed to enhance smart agent simulation for autonomous driving.
+This work introduced BehaviorGPT, a fully autoregressive architecture designed to enhance smart agent simulation for autonomous driving.
 By applying homogeneous Transformer blocks to entire trajectory snippets and utilizing relative spacetime representations, BehaviorGPT simplifies the modeling process and maximizes data utilization.
-To enable high-level understanding and long-range interaction reasoning in space and time, we developed the Next-Patch Prediction Paradigm, which requires the model to predict patches of trajectories instead of short segments.
-Our results on the Waymo Open Sim Agents Challenge demonstrate that BehaviorGPT, with just 3 million parameters, achieves superior performance, highlighting its potential to improve the realism and scalability of autonomous vehicle simulations with significantly fewer parameters.
+To enable high-level understanding and long-range interaction reasoning in space and time, we developed the Next-Patch Prediction Paradigm, which tasks models with generating trajectory patches instead of single-step states.
+Experimental results on the Waymo Open Sim Agents Challenge demonstrate that BehaviorGPT achieves outstanding performance with merely 3M model parameters, highlighting its potential to further improve the realism of agent simulation with more data and computation.
+Limitations.
+First, BehaviorGPT is currently inferior in kinematics-related performance, which can be enhanced by incorporating a kinematic model, e.g., the bicycle model.
+Second, the current version of BehaviorGPT does not support controlling agent behavior with specific prompts such as language and goal points.
+However, achieving controllable generation should be trivial given a powerful base model.
+Finally, we have not verified whether BehaviorGPT will facilitate the development of motion planning, which we leave as future work.
+Acknowledgement
+This project is supported by a grant from Hong Kong Research Grant Council under GRF project 11216323 and CRF C1042-23G.

### v3

@@
 Patch Size | Replan Frequency | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
-| 1 | 10 Hz | 2.3752 | 0.6783 | 0.2559 | 0.2088 | 0.4022 | 0.5094 | 0.3201 | 0.9002 | 0.8015 | 0.6149 | 0.8432 |
-| 5 | 2 Hz | 1.5599 | 0.7273 | 0.3543 | 0.3218 | 0.4623 | 0.5435 | 0.3768 | 0.9181 | 0.8339 | 0.6564 | 0.9077 |
-| 10 | 1 Hz | 1.5203 | 0.7335 | 0.3517 | 0.3023 | 0.4734 | 0.5432 | 0.3797 | 0.9358 | 0.8329 | 0.6645 | 0.9132 |
+1 | 10 Hz | 2.3752 | 0.6783 | 0.2559 | 0.2088 | 0.4022 | 0.5094 | 0.3201 | 0.9002 | 0.8015 | 0.6149 | 0.8432 |
+5 | 2 Hz | 1.5599 | 0.7273 | 0.3543 | 0.3218 | 0.4623 | 0.5435 | 0.3768 | 0.9181 | 0.8339 | 0.6564 | 0.9077 |
+10 | 1 Hz | 1.5203 | 0.7335 | 0.3517 | 0.3023 | 0.4734 | 0.5432 | 0.3797 | 0.9358 | 0.8329 | 0.6645 | 0.9132 |
 Impact of replan frequency.
@@
 Patch Size | Replan Frequency | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
-| 10 | 1 Hz | 1.5405 | 0.7414 | 0.3553 | 0.3153 | 0.4695 | 0.5303 | 0.3772 | 0.9520 | 0.8285 | 0.6664 | 0.9308 |
-| 10 | 2 Hz | 1.4147 | 0.7473 | 0.3615 | 0.3365 | 0.4806 | 0.5544 | 0.3834 | 0.9537 | 0.8308 | 0.6702 | 0.9349 |
-| 10 | 5 Hz | 1.5693 | 0.7342 | 0.3430 | 0.3472 | 0.4663 | 0.5673 | 0.3722 | 0.9429 | 0.8253 | 0.6534 | 0.9089 |
+10 | 1 Hz | 1.5405 | 0.7414 | 0.3553 | 0.3153 | 0.4695 | 0.5303 | 0.3772 | 0.9520 | 0.8285 | 0.6664 | 0.9308 |
+10 | 2 Hz | 1.4147 | 0.7473 | 0.3615 | 0.3365 | 0.4806 | 0.5544 | 0.3834 | 0.9537 | 0.8308 | 0.6702 | 0.9349 |
+10 | 5 Hz | 1.5693 | 0.7342 | 0.3430 | 0.3472 | 0.4663 | 0.5673 | 0.3722 | 0.9429 | 0.8253 | 0.6534 | 0.9089 |
 Agent-Agent Self-Attention | minADE () | REALISM () | DIST TO OBJ () | COLLISION () | TTC () |
-| 2.1489 | 0.6659 | 0.3539 | 0.6987 | 0.8070 | |
+2.1489 | 0.6659 | 0.3539 | 0.6987 | 0.8070 | |
 ✓ | 1.6247 | 0.7349 | 0.3783 | 0.9409 | 0.8320 |
@@
 Model Depth | #Param | minADE () | REALISM () |
-| 2 | 3M | 1.6247 | 0.7349 |
-| 3 | 4M | 1.5381 | 0.7387 |
-| 4 | 5M | 1.4881 | 0.7396 |
+2 | 3M | 1.6247 | 0.7349 |
+3 | 4M | 1.5381 | 0.7387 |
+4 | 5M | 1.4881 | 0.7396 |
 Model Width | #Param | minADE () | REALISM () |
-| 64 | 800K | 1.9637 | 0.7251 |
-| 128 | 3M | 1.6247 | 0.7349 |
-| 192 | 7M | 1.4993 | 0.7382 |
+64 | 800K | 1.9637 | 0.7251 |
+128 | 3M | 1.6247 | 0.7349 |
+192 | 7M | 1.4993 | 0.7382 |
 Training | Inference | minADE () | REALISM () | LINEAR SPEED () | LINEAR ACCEL () | ANG SPEED () | ANG ACCEL () | DIST TO OBJ () | COLLISION () | TTC () | DIST TO ROAD EDGE () | OFFROAD () |
-| 9.1 sec | 9.1 sec | 1.6247 | 0.7349 | 0.3546 | 0.3105 | 0.4689 | 0.5363 | 0.3783 | 0.9409 | 0.8320 | 0.6605 | 0.9163 |
-| 5.0 sec | 9.1 sec | 1.6294 | 0.7333 | 0.3565 | 0.3471 | 0.4613 | 0.5293 | 0.3813 | 0.9375 | 0.8273 | 0.6585 | 0.9100 |
+9.1 sec | 9.1 sec | 1.6247 | 0.7349 | 0.3546 | 0.3105 | 0.4689 | 0.5363 | 0.3783 | 0.9409 | 0.8320 | 0.6605 | 0.9163 |
+5.0 sec | 9.1 sec | 1.6294 | 0.7333 | 0.3565 | 0.3471 | 0.4613 | 0.5293 | 0.3813 | 0.9375 | 0.8273 | 0.6585 | 0.9100 |
 Scaling with data.